OpenStudy (anonymous):

James bought 12 cookies and ate 3 of them. He ate ____% of the cookies. THE RIGHT ANSWER GETS A MEDAL AND I WILL BECOME A FAN!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

Since 12 cookies equals 100% of the cookies. The part is 3 cookies. So set up a proportion to solve. \[\frac{ part }{ whole }= \frac{ ? }{ 100 }\] \[\frac{ 3 }{12 }=\frac{ ? }{ 100 }\] To solve, cross multiply (3*100) and divide by 12. 3*100=300 300 divided by 12 = 25 So he ate 25% of the cookies.